Hello! Thank you for your time. I've been trying to solve this for a week now.
Solve: X=51n(7+3y)for y.



The answer is supposed to be:
x
-
e5-7
____
3
This is the same answer without the Division bars
x
e5-7
3







Found 2 solutions by mukhopadhyay, Prithwis:
Answer by mukhopadhyay(490)

x = 51n(7+3y)
=> ln(7+3y) = x/5;
=> (7+3y) = e^(x/5);
=> 3y = e^(x/5) - 7
=> y = 1/3[e^(x/5) - 7]

Answer by Prithwis(166)

x = 51n(7+3y)
=> ln(7+3y) = x/5;
=> (7+3y) = e^(x/5);
=> 3y = e^(x/5) - 7
=> y = 1/3[e^(x/5) - 7]